# =COLI-manual3.qry Cross-lingual Cross-temporal Summarization: Dataset, Models, Evaluation

### v1

@@ Abstract
While summarization has been extensively researched in natural language processing (NLP), cross-lingual cross-temporal summarization (CLCTS) is a largely unexplored area that has the potential to improve cross-cultural accessibility, information sharing, and understanding. This paper comprehensively addresses the CLCTS task, including dataset creation, modeling, and evaluation. We build the first CLCTS corpus, leveraging historical fictive texts and Wikipedia summaries in English and German, and examine the effectiveness of popular transformer end-to-end models with different intermediate task finetuning tasks. Additionally, we explore the potential of ChatGPT for CLCTS as a summarizer and an evaluator. Overall, we report evaluations from humans, ChatGPT, and several recent automatic evaluation metrics where we find our intermediate task finetuned end-to-end models generate bad to moderate quality summaries; ChatGPT as a summarizer (without any finetuning) provides moderate to good quality outputs and as an evaluator correlates moderately with human evaluations though it is prone to giving lower scores. ChatGPT also seems to be very adept at normalizing historical text. We finally test ChatGPT in a scenario with adversarially attacked and unseen source documents and find that ChatGPT is better at omission and entity swap than negating against its prior knowledge.
@@ 1 Introduction
Summarization is a key task in natural language processing (NLP), especially in an age of information overload. The classical approach for summarization is to summarize documents (e.g., news reports) in the same language Zhang et al. (2020a); Liu et al. (2022); Ravaut, Joty, and Chen (2022) or, less prominently, in different languages than the original source (cross-lingual summarization; CLS) Ladhak et al. (2020); Cao, Liu, and Wan (2020); Bai, Gao, and Huang (2021); Liang et al. (2022). Tasks such as the summarization of historical texts to modern languages (cross-temporal summarization; CTS) are scarcely explored, in contrast. Such summarizers are beneficial not only for historical researchers and students but also for laypeople with an interest in historical information sources.
In this work, we go one step further and explore cross-lingual cross-temporal summarization (CLCTS), in which the goal is to summarize a historical document in a different modern language. See Table 1 for an example. Such cross-lingual and cross-temporal summarizers would potentially tremendously facilitate cross-cultural accessability, information sharing, and understanding.
Title |
An Inhabitant of Carcosa |
Author |
Ambrose Bierce |
Year |
1886 |
Source |
The Collected Works of Ambrose Bierce, Volume 3 |
Text |
@@
“For there be divers sorts of death—some wherein the body remaineth; and in some it vanisheth quite away with the spirit … In one kind of death the spirit also dieth, and this it hath been known to do while yet the body was in vigor for many years. […]” |
Summary (German) |
@@
“Ein Mann aus der Stadt Carcosa, der über die Worte des Philosophen Hali über die Natur des Todes nachdenkt, wandert durch eine ihm unbekannte Wildnis. […]” |
@@
CLCTS is difficult and challenging for the following reasons: (i) languages change over time along multiple dimensions, including syntax Juzek, Krielke, and Teich (2020); Lei and Wen (2020), semantics Hamilton, Leskovec, and Jurafsky (2016); Giulianelli, Del Tredici, and Fernández (2020), lexical choice and morphology Gibson et al. (2019); Joseph (2017). (ii) Historical documents are often longer, not only because of tendencies of simplification/text length reduction over time Lewis (1894); Sherman (1893); Rudnicka (2018), but also because of the genre of historical text, which predominantly includes literary work Wang and Liu (2017); Zhu, Liu, and Pang (2022). (iii) The process can be thought of as consisting of several subtasks, previously considered independent in NLP research: machine translation (MT), summarization and historical (spelling) normalization Bollmann and Søgaard (2016); Eger, vor der Brück, and Mehler (2016).
@@
In this work, we consider the CLCTS problem comprehensively, including dataset creation, modeling, and evaluation.111Code+data: https://github.com/zhangr2021/CLCTS We start with building the first CLCTS dataset. Our corpus contains stories and plays (as prime artefacts of historical cultures), which requires the models to handle long documents; we assemble modern summaries by querying Wikipedia.
@@
In terms of modeling, we consider three approaches: (1) extractive summarizers Gu, Ash, and Hahnloser (2022) whose advantage is that they can in principle deal with unlimited input document lengths but which are limited in that they can only copy out sentences from the original source (which is a problem, e.g., when the source is first-person narrative but the summary should be third-person perspective). (2) Abstractive summarizers can avoid the issue above but typically have severe limitations in input lengths. Even models exclusively designed for long document summarization by utilizing efficient variants of attention mechanism such as Longformer Encoder-Decoder (LED) Beltagy, Peters, and Cohan (2020) and Bigbird Zaheer et al. (2020) limit the max input tokens to 16,384 and 4,096, respectively. (3) GPT-3 Brown et al. (2020) and its extensions such as ChatGPT are also strong candidates for our task Goyal, Li, and Durrett (2022) as they have so spectacularly impacted research in a multitude of scenarios Leiter et al. (2023), including summarization Yang et al. (2023); Wang et al. (2023); Zhang, Liu, and Zhang (2023). Consequently, we also include ChatGPT in our analysis.
@@
In terms of evaluation, we consider three evaluation methods: (1) Automatic evaluation, where we report multiple evaluation metrics additional to the commonly used variants of ROUGE Lin (2004) in summarization tasks Liang et al. (2022); Bai, Gao, and Huang (2021); Cao, Liu, and Wan (2020); Peng et al. (2021) since the performance of individual metrics may vary across datasets, challenging their reliability Bhandari et al. (2020); Fabbri et al. (2021). (2) Human evaluation, which reflects the actual quality of summaries according to human judgements and functions as a source of reliability measurement for automatic evaluation metrics. (3) ChatGPT evaluation, where we examine the potential of ChatGPT as an alternative to human annotators given the same instruction considering the high cost of human evaluation Gao et al. (2023) and issues of reproducibility Belz et al. (2021); Belz, Thomson, and Reiter (2023); Chen, Belouadi, and Eger (2022).
Our contributions are:
@@
To our best knowledge, we build the first CLCTS corpus, leveraging historical fictive texts and Wikipedia summaries.
@@
We provide both human, ChatGPT and automatic evaluation for the task, showing that our intermediate task finetuned e2e models generate bad to moderate quality summaries; ChatGPT as a summarizer provides moderate to good quality outputs (without any finetuning); ChatGPT as an evaluator correlates with human evaluations to a moderate level though it is prone to lower scores.
@@ 2 Related work
Since we focus on various summarization instead of translation and long documents instead of paragraphs,
the translation dataset is not within the scope of our current study.
@@
attention (input limits 4096 and 16384) and PEGASUS with BigBird attention (4096).333The very recently proposed retrieval-based approach unlimiformer Bertsch et al. (2023) extends the input limitations of the aforementioned pretrained encoder-decoder transformers to practically unlimited input sequences by offloading the cross-attention computation.
We build our model with mBART
Tang et al. (2020) pretrained for multilingual MT tasks and combine with LED attention to enable efficient processing for long sequences.
Additionally, we build another retrieve-then-summarize pipeline
where we retrieve sentences with an extractive summarizer Gu, Ash, and Hahnloser (2022) and then summarize cross-lingually using ChatGPT.
ChatGPT for summarization Despite ChatGPT’s young age, there is already a wealth of research that explores it for various NLP tasks, including summarization Yang et al. (2023) and summarization evaluation Chiang and yi Lee (2023); Gao et al. (2023). Disagreement exists concerning ChatGPT’s performance.
@@
in an interactive manner. It is also worth noticing that all works mentioned above utilize small samples of 50 to 100 random samples (out of 3000 to 60k observations) from each dataset in their experiments.
Summarization evaluation metrics Depending on the availability of reference summaries, one can utilize reference-based metrics (focusing on overlap measure Lin (2004); Popović (2015), embedding similarity Zhang et al. (2020b); Zhao et al. (2019), discourse coherence Zhao, Strube, and Eger (2023), natural language inference (NLI) Chen and Eger (2023) and text generation task such as question answering (QA) Yuan, Neubig, and Liu (2021); Wang, Cho, and Lewis (2020)) or reference-free metrics Gao, Zhao, and Eger (2020); Chen, Li, and King (2021); Liu, Jia, and Zhu (2022); Laban et al. (2022); Belouadi and Eger (2023).
@@
To address this concern, we leverage multiple automatic evaluation metrics, especially more recent ones e.g., MENLI and DiscoScore, as well as human evaluation. Since we focus on long documents,
we do not explore reference-free metrics, due to the vast mismatch in length between source texts and summaries.
@@ 3.1 CLCTS Corpus
Manually creating summaries is very time-consuming and requires expert knowledge Peng et al. (2021). Thus, we use fairy tales, short stories and plays where summaries are available in Wikipedia articles. In each language direction, English and German, the dataset contains summary pairs from historical documents in one language to modern summaries in the other language.
Dataset collection We collect historical documents mainly from three different sources. The links to the sources are collected in Appendix 9.1.
@@
Our workflow for building the corpus consists of two stages: (1) meta-information collection & historical text extraction and (2) summary collection & translation. In stage (1), we collect meta-information of historical documents using the “Beautifulsoup” library. We extract information such as titles, web links, authors, year, and documents together with information from Wikipedia lists of German fairy tales and English short stories. For documents presented in several sources, the oldest version or the version with historical spelling is selected. For stage (2), we utilize the collected information to match the historical documents with their corresponding summaries from Wikipedia. The historical spelling of titles is converted to modern spelling to match the titles in Wikipedia. Figure 1 illustrates the process of data collection. If no original human correspondent summaries are found in the other language, we translate the summaries into the target language using DeepL (we translate 144 out of 328 German summaries to English and 108 out of 289 English summaries to German).555DeepL API: https://www.deepl.com/pro-api An example of our collected dataset is shown in Table 1.
In addition to the summarization datasets, we create a historical German-English translation dataset with 201 translation pairs. This dataset is later utilized in intermediate task finetuning. We retrieve the German documents from the DTA corpus and the historical English from Wikisource. The documents are translated by Margaret Raine Hunt (1831–1912). An example of the translation dataset is in Table 13 (appendix).
@@ 3.2 External datasets
To enrich our finetuning data resources, we utilize external summarization datasets: (i) HISTSUMM Peng et al. (2021) where we also translate the summaries into English using DeepL to enable cross-lingual finetuning. (ii) We further use Wikilingua for MLS and CLS tasks Ladhak et al. (2020). Wikilingua is a multilingual summarization dataset which contains 18 languages from WikiHow (an online wiki-style publication featuring how-to articles), including English and German. The English dataset contains 141,457 document-summary pairs. For finetuning our models, we randomly select 60% of the pairs (85,874) and we use all of the German dataset (58,341). Wikilingua German entries contain the URL and the section name of the English articles. Thus, cross-lingual summary pairs can be retrieved using this information and we match 20,103 cross-lingual summary pairs in both directions. (iii) We also utilize the CNN/Daily Mail dataset. This is a monolingual English summarization dataset containing 311,971 document-summary pairs, including news articles written by CNN and Daily Mail journalists, and summaries from the highlight of the article written by the article author. We experiment with a subset of 86,133 text-summary pairs. (iv) We include MLSUM (Scialom et al., 2020), which is a multilingual summarization dataset containing over 1.5 million document-summary pairs collected from online newspapers such as Süddeutsche Zeitung (German). We use a subset of 66,226 pairs from the German dataset.
@@ 3.3 Dataset Statistics
CLCTS hEn-De | 289 | 9,643.2 | 413.9 | 16.0 | 19.2 | 23.3 | fiction | CLS, CTS |
CLCTS hDe-En | 328 | 1,416.8 | 324.1 | 25.4 | 18.6 | 4.4 | fiction | CLS, CTS |
CLCTS hEn-En | 289 | 9,643.2 | 423.6 | 16.0 | 20.2 | 22.8 | fiction | MLS, CTS |
CLCTS hDe-De | 328 | 1,416.8 | 217.3 | 25.4 | 16.6 | 6.5 | fiction | MLS, CTS |
@@
Much fewer instances in Historical datasets. The CLCTS hEn-De dataset consists of 289 historical short stories and plays, and the CLCTS hDe-En dataset consists of 328 historical fairy tales. The size of HISTSUMM dataset is even smaller, with only 100 historical news articles. In contrast, modern summarization datasets contain several thousand summarization pairs, which is more than 100 times the size of historical datasets. The largest dataset CNN/Daily Mail contains over 300k summarization pairs.
Longer documents and summaries for CLCTS corpus. The CLCTS hEn-De dataset has more than 9643 tokens per document and 413 tokens per summary on average and CLCTS hDe-En has 1416 per document and 324 per summary. For external datasets, the average length of documents is below 800 tokens per document and below 60 tokens per summary. This is because the CLCTS corpus includes historical documents from the genre fiction, which are naturally longer than news and WikiHow (how-to guide) due to genre differences (Rudnicka, 2018) and higher verbosity of more historical languages.
Larger information loss during summarization for news genre and CLCTS hEn-De measured by compression ratio. News documents have a higher compression ratio (the mean length of document divided by the mean length of summary) compared to documents from other genres, except for CLCTS hEn-De, which has the highest compression ratio among all datasets. The CLCTS hEn-De dataset contains short stories and plays which are much longer than fairytales in CLCTS hDe-En and documents from all other genres as discussed above. For the CLCTS corpus, the compression ratio differs substantially because all summaries of the CLCTS corpus are collected from Wikipedia with the average length of summaries ranging from 200 to 400 tokens regardless of the original document length.

### v3

# Cross-lingual Cross-temporal Summarization: Dataset, Models, Evaluation

@@ Abstract
While summarization has been extensively researched in natural language processing (NLP), cross-lingual cross-temporal summarization (CLCTS) is a largely unexplored area that has the potential to improve cross-cultural accessibility and understanding. This paper comprehensively addresses the CLCTS task, including dataset creation, modeling, and evaluation. We (1) build the first CLCTS corpus with 328 instances for hDe-En (extended version with 455 instances) and 289 for hEn-De (extended version with 501 instances), leveraging historical fiction texts and Wikipedia summaries in English and German; (2) examine the effectiveness of popular transformer end-to-end models with different intermediate finetuning tasks; (3) explore the potential of GPT-3.5 as a summarizer; (4) report evaluations from humans, GPT-4, and several recent automatic evaluation metrics. Our results indicate that intermediate task finetuned end-to-end models generate bad to moderate quality summaries while GPT-3.5, as a zero-shot summarizer, provides moderate to good quality outputs. GPT-3.5 also seems very adept at normalizing historical text. To assess data contamination in GPT-3.5, we design an adversarial attack scheme in which we find that GPT-3.5 performs slightly worse for unseen source documents compared to seen documents. Moreover, it sometimes hallucinates when the source sentences are inverted against its prior knowledge with a summarization accuracy of 0.67 for plot omission, 0.71 for entity swap, and 0.53 for plot negation. Overall, our regression results of model performances suggest that longer, older, and more complex source texts (all of which are more characteristic for historical language variants) are harder to summarize for all models, indicating the difficulty of the CLCTS task. Regarding evaluation, we observe that both GPT-4 and BERTScore correlate moderately with human evaluations, implicating great potential for future improvement.
@@ 1 Introduction
Summarization is a key task in natural language processing (NLP), especially in an age of information overload. The classical approach for summarization is to summarize documents (e.g., news reports) in the same language Zhang et al. (2020a); Liu et al. (2022); Ravaut, Joty, and Chen (2022) or, less prominently, in different languages than the source (cross-lingual summarization; CLS) Ladhak et al. (2020); Cao, Liu, and Wan (2020); Bai, Gao, and Huang (2021); Liang et al. (2022). Tasks such as the summarization of historical texts to modern languages (cross-temporal summarization; CTS) are scarcely explored, in contrast. Such summarizers are beneficial not only for historical researchers and students but also for laypeople with an interest in historical information sources.
In this work, we go one step further and explore cross-lingual cross-temporal summarization (CLCTS), in which the goal is to summarize a historical document in a different modern language. See Table 1 for an example. Such cross-lingual and cross-temporal summarizers would potentially tremendously facilitate cross-cultural accessibility, information sharing, and understanding. CLCTS is a valuable area of research that expands upon the recently popular task of literary translation with summarization. Both tasks share cultural importance and present unique complexities due to the intricate nature of the creative work involved Karpinska and Iyyer (2023); Thai et al. (2022). CLCTS is not limited to literary works alone: it can also assist in summarizing non-literature such as historical newspapers, reports, textbooks, etc. For example, CLCTS models can assist in creating (missing) Wikipedia pages for historical works. While the translation of literature may require language that is faithful to the original source for a ‘close reader’, CLCTS could potentially be better suited for a ‘distant reader’, and thus may require less sophisticated output text. Moreover, modern summaries can facilitate the classification of literary works, for example, genre classification Saraswat and Srishti (2022); Agarwal, Vijay et al. (2021). Firstly, summaries are more condensed than (historical) source texts and, therefore, are much easier to model. Secondly, modern texts are better suited for models that operate with modern data such as SBERT Reimers and Gurevych (2020).
Title | An Inhabitant of Carcosa |
Author | Ambrose Bierce |
Year | 1886 |
Source | The Collected Works of Ambrose Bierce, Volume 3 |
@@
Text | “For there be divers sorts of death—some wherein the body remaineth; and in some it vanisheth quite away with the spirit … In one kind of death the spirit also dieth, and this it hath been known to do while yet the body was in vigor for many years. […]” |
@@
Summary (German) | “Ein Mann aus der Stadt Carcosa, der über die Worte des Philosophen Hali über die Natur des Todes nachdenkt, wandert durch eine ihm unbekannte Wildnis. […]” |
@@
CLCTS is difficult and challenging for the following reasons: (i) languages change over time along multiple dimensions, including syntax Juzek, Krielke, and Teich (2020); Lei and Wen (2020), semantics Hamilton, Leskovec, and Jurafsky (2016); Giulianelli, Del Tredici, and Fernández (2020), lexical choice and morphology Gibson et al. (2019); Joseph (2017). (ii) Historical documents are often longer, not only because of tendencies of simplification/text length reduction over time Lewis (1894); Sherman (1893); Rudnicka (2018), but also because of the genre of historical text, which predominantly includes literary work Wang and Liu (2017); Zhu, Liu, and Pang (2022). (iii) The process can be thought of as consisting of several subtasks, previously considered independently in NLP research: machine translation (MT), summarization, and historical (spelling) normalization Bollmann and Søgaard (2016); Eger, vor der Brück, and Mehler (2016).
@@
In this work, we consider the CLCTS problem comprehensively, including dataset creation, modeling, and evaluation.111Our code and data are available at https://github.com/zhangr2021/CLCTS. We start with building the first CLCTS dataset. Our corpus contains stories and plays (as prime artefacts of historical cultures), which requires the models to handle long documents; we assemble modern summaries by querying Wikipedia.
@@
In terms of modeling, we consider three approaches: (1) extractive summarizers Gu, Ash, and Hahnloser (2022) whose advantage is that they can in principle deal with unlimited input document lengths but which are limited in that they can only copy out sentences from the original source (which is a problem, e.g., when the source is first-person narrative but the summary should be third-person perspective). (2) Abstractive summarizers can avoid the issue above but typically have severe limitations in input lengths. Even models exclusively designed for long document summarization by utilizing efficient variants of attention mechanism such as Longformer Encoder-Decoder (LED) Beltagy, Peters, and Cohan (2020) and Bigbird Zaheer et al. (2020) limit the max input tokens to 16,384 and 4,096, respectively. (3) GPT-3 Brown et al. (2020) and its extensions such as ChatGPT are also strong candidates for our task Goyal, Li, and Durrett (2022a) as they have so spectacularly impacted research in a multitude of scenarios Leiter et al. (2024), including summarization Yang et al. (2023); Wang et al. (2023); Zhang, Liu, and Zhang (2023). Consequently, we also include ChatGPT (GPT-3.5 and GPT-4) in our analysis.
@@
In terms of evaluation, we consider three evaluation methods: (1) automatic evaluation, where we report multiple evaluation metrics additional to the commonly used variants of ROUGE Lin (2004) in summarization tasks Liang et al. (2022); Bai, Gao, and Huang (2021); Cao, Liu, and Wan (2020); Peng et al. (2021) since the performance of individual metrics may vary across datasets, challenging their reliability Bhandari et al. (2020); Fabbri et al. (2021); (2) human evaluation, which reflects the actual quality of summaries according to human judgments and functions as a source of reliability measurement for automatic evaluation metrics; (3) ChatGPT evaluation, where we examine the potential of ChatGPT as an alternative to human annotators given the same instruction considering the high cost of human evaluation Gao et al. (2023) and issues of reproducibility Belz et al. (2021); Belz, Thomson, and Reiter (2023); Chen, Belouadi, and Eger (2022). Our contributions are:
@@
To our best knowledge, we build the first CLCTS corpus, leveraging historical fiction texts and Wikipedia summaries.
@@
We provide human, GPT-4, and automatic evaluation for the task, showing that our intermediate task finetuned e2e models generate bad to moderate quality summaries; GPT-3.5 as a summarizer provides moderate to good quality outputs (without any finetuning); GPT-4 as an evaluator correlates with human evaluations to a moderate level.
@@ 2 Related work
Historical text normalization Historical text normalization, the process of mapping non-standard word tokens to their modern standard forms, is an important technique for analyzing historical texts. Model designs include (1) corpus-based lexical substitution as an effective component for normalization systems Rayson, Archer, and Smith (2005); Baron and Rayson (2008); Bollmann (2012); (2) rule-based models, such as phonological rules Porta, Sancho, and Gómez (2013) and edit distance measures Adesam, Ahlberg, and Bouma (2012); Bollmann (2012); (3) statistical models which aim to maximize the probability of contemporary word form given its historical form Pettersson (2016); (4) neural network (NN) models which utilize NN architectures such as encoder-decoder LSTMs Bollmann, Bingel, and Søgaard (2017); Robertson and Goldwater (2018) or RNNs Makarov and Clematide (2020). Worth noting is that nearly all existing historical normalizers operate on words in isolation, except for the semi-supervised contextual normalizer from Makarov and Clematide (2020). In our study, we find that ChatGPT (GPT-3.5) has the potential to serve as a (much better) context-aware historical normalizer.
@@
Since we focus on various summarizations instead of translations and long documents instead of paragraphs, the translation dataset is not within the scope of our current study.
@@
We build our model with mBART Tang et al. (2020) pretrained for multilingual MT tasks and combine with LED attention to enable efficient processing for long sequences. Additionally, we build another retrieve-then-summarize pipeline where we retrieve sentences with an extractive summarizer Gu, Ash, and Hahnloser (2022) and then summarize cross-lingually using ChatGPT.
ChatGPT for summarization Despite ChatGPT’s young age, there is already a wealth of research that explores it for various NLP tasks, including summarization Yang et al. (2023) and summarization evaluation Chiang and Lee (2023); Gao et al. (2023). Disagreement exists concerning ChatGPT’s performance.
@@
in an interactive manner. It is also worth noticing that all works mentioned above utilize small samples of 50 to 100 random samples (out of 3,000 to 60k observations) from each dataset in their experiments.
@@ 3.1 CLCTS Corpus
Manually creating summaries is very time-consuming and requires expert knowledge Peng et al. (2021). Thus, we use fairy tales, short stories, and plays where summaries are available in Wikipedia articles. In each language direction, English and German, the dataset contains summary pairs from historical documents in one language to modern summaries in the other language.
Dataset collection We collect historical documents mainly from three different sources.555Wikisource and Project Gutenberg contain both German and English sources, while Deutsches Textarchiv contains German sources only. The links to the sources are collected in Section 9.1 (appendix).
@@
Our workflow for building the corpus consists of two stages: (1) meta-information collection & historical text extraction and (2) summary collection & translation. In stage (1), we collect meta-information of historical documents using the “Beautifulsoup” library. We extract information such as titles, web links, authors, years, and documents together with information from Wikipedia lists of German fairy tales and English short stories. For documents presented in several sources, the oldest version or the version with historical spelling is selected. For stage (2), we utilize the collected information to match the historical documents with their corresponding summaries from Wikipedia. The historical spelling of titles is converted to modern spelling to match the titles in Wikipedia. Figure 1 illustrates the process of data collection. If no original human correspondent summaries are found in the other language, we translate the summaries into the target language using DeepL (we translate 144 out of 328 German summaries to English and 108 out of 289 English summaries to German).666DeepL API: https://www.deepl.com/pro-api An example of our collected dataset is shown in Table 1.
In addition to the summarization datasets, we create a historical German-English translation dataset with 201 translation pairs. This dataset is later utilized in intermediate task finetuning. We retrieve the German documents from the DTA corpus and the historical English from Wikisource. The documents are translated by Margaret Raine Hunt (1831–1912). An example of the translation dataset is given in Table 17 (appendix).
@@ 3.2 External datasets
To enrich our finetuning data resources, we utilize external summarization datasets: (1) HISTSUMM Peng et al. (2021) where we also translate the summaries into English using DeepL to enable cross-lingual finetuning. (2) We further use Wikilingua for MLS and CLS tasks Ladhak et al. (2020). Wikilingua is a multilingual summarization dataset that contains 18 languages from WikiHow (an online Wiki-style publication featuring how-to articles), including English and German. The English dataset contains 141,457 document-summary pairs. For finetuning our models, we randomly select 60% of the pairs (85,874) and we use all of the German dataset (58,341). Wikilingua German entries contain the URL and the section name of the English articles. Thus, cross-lingual summary pairs can be retrieved using this information and we match 20,103 cross-lingual summary pairs in both directions. (3) We also utilize the CNN/Daily Mail dataset. This is a monolingual English summarization dataset containing 311,971 document-summary pairs, including news articles written by CNN and Daily Mail journalists, and summaries from the highlight of the article written by the article author. We experiment with a subset of 86,133 text-summary pairs. (4) We include MLSUM (Scialom et al., 2020), which is a multilingual summarization dataset containing over 1.5 million document-summary pairs collected from online newspapers such as Süddeutsche Zeitung (German). We use a subset of 66,226 pairs from the German dataset.
@@ 3.3 Dataset Statistics
CLCTS hEn-De | 289 | 9,643.3 | 414.0 | 16.3 | 19.2 | 23.3 | fiction | CLS, CTS |
CLCTS hDe-En | 328 | 1,398.4 | 324.3 | 25.4 | 18.6 | 4.3 | fiction | CLS, CTS |
CLCTS hEn-En | 276 | 9,847.2 | 488.1 | 16.0 | 20.2 | 20.2 | fiction | MLS, CTS |
CLCTS hDe-De | 328 | 1,398.4 | 217.4 | 25.4 | 16.6 | 6.4 | fiction | MLS, CTS |
@@
Much fewer instances in Historical datasets. The CLCTS hEn-De dataset consists of 289 historical short stories and plays, and the CLCTS hDe-En dataset consists of 328 historical fairy tales. 777We limit the maximum length of the document to 16k. Since we collect summaries from Wikipedia, the total number of matched document-summary pairs is limited. Stories with less popularity are less likely to be found on Wikipedia than well-known ones. The size of HISTSUMM dataset is even smaller, with only 100 historical news articles. In contrast, modern summarization datasets contain several thousand summarization pairs, which is more than 100 times the size of historical datasets. The largest dataset CNN/Daily Mail contains over 300k summarization pairs.
Longer documents and summaries for CLCTS corpus. The CLCTS hEn-De dataset has more than 9,643 tokens per document and 414 tokens per summary on average and CLCTS hDe-En has 1,398 per document and 324 per summary. For external datasets, the average length is below 800 tokens per document and below 60 tokens per summary. This is because the CLCTS corpus includes historical documents from the genre fiction, which are naturally longer than news and WikiHow (how-to guide) due to genre differences (Rudnicka, 2018) and higher verbosity of more historical languages.
Larger information loss during summarization for news genre and CLCTS hEn-De measured by compression ratio. News documents have a higher compression ratio (the mean length of document divided by the mean length of summary) compared to documents from other genres, except for CLCTS hEn-De, which has the highest compression ratio among all datasets. The CLCTS hEn-De dataset contains short stories and plays that are much longer than fairytales in CLCTS hDe-En and documents from all other genres as discussed above. For the CLCTS corpus, the compression ratio differs substantially because all summaries of the CLCTS corpus are collected from Wikipedia with the average length of summaries ranging from 200 to 400 tokens regardless of the original document length.